# Social Good or Scientific Curiosity? Uncovering the Research Framing Behind NLP Artefacts

### v1

@@ Abstract
Clarifying the research framing of NLP artefacts (e.g., models, datasets, etc.) is crucial to aligning research with practical applications. Recent studies manually analyzed NLP research across domains, showing that few papers explicitly identify key stakeholders, intended uses, or appropriate contexts. In this work, we propose to automate this analysis, developing a three-component system that infers research framings by first extracting key elements (means, ends, stakeholders), then linking them through interpretable rules and contextual reasoning. We evaluate our approach on two domains: automated fact-checking using an existing dataset, and hate speech detection for which we annotate a new dataset [^1] Code and annotations available at our GitHub repository. —achieving consistent improvements over strong LLM baselines. Finally, we apply our system to recent automated fact-checking papers and uncover three notable trends: a rise in vague or underspecified research goals, increased emphasis on scientific exploration over application, and a shift toward supporting human fact-checkers rather than pursuing full automation.
@@ 1 Introduction
As NLP systems see wider real-world deployment, concerns over misuse and unintended consequences have intensified focus on responsible AI practices—highlighting the need for ethical integration, risk mitigation, and attention to societal impact (Noble, 2018; Bender et al., 2021; Birhane et al., 2022). A critical but often overlooked aspect of responsible NLP research is research framing—how a paper positions the purpose and value of its work by connecting what it does (means), why it does it (ends), and for whom it is relevant (stakeholders). When poorly articulated, framings can obscure a system’s scope, hinder accountability, and increase the risk of misuse. Figure 1 illustrates this point by comparing two NLP papers: one clearly intends to support professional fact-checkers with specified means, ends, and users; the other, on hate speech detection, lacks any mention of who would use the model’s outputs. Such omissions risk misinterpretation—e.g., framing a label-only classification model as an automated decision tool overlooks the need for context and explanation, risking misuse in high-stakes settings.
@@
Figure 1: Side-by-side examples of research framings. The fact-checking paper (left, Konstantinovskiy et al. (2021)) presents a clear assisted external fact-checking framing with specified users, goals, and methods. The hate speech paper (right, Ousidhoum et al. (2019)) has a vague framing due to unspecified model users.
This concern has motivated several analyses of how NLP research is framed. Schlichtkrull et al. (2023) first formalized the notion of research framings (which they termed “epistemic narratives”) and introduced epistemic elements—textual signals that define a paper’s framing by specifying components such as means, ends and stakeholders. Analyzing 100 automated fact-checking papers, they found frequent misalignments between goals and methods, vague or missing stakeholder definitions, and weak links between these elements. Separately, Liu et al. (2023) reviewed 333 summarization papers and found that fewer than 15% meaningfully addressed responsible AI concerns.
@@
    Epistemic element extraction, i.e., identifying epistemic elements from introductory text.
@@
We evaluate our approach (§5) in two domains: automated fact-checking (AFC) and hate speech detection (HS). For AFC, we use the dataset released by Schlichtkrull et al. (2023); for HS, we manually annotate 49 highly cited papers (§4) with epistemic elements, which we use to infer their research framings. Our approach outperforms a strong LLM baseline in both domains, offering more accurate research framing predictions. It also achieves high precision in identifying vague framings with misaligned or incomplete elements.
Finally, we apply our system (§6) to recent AFC papers to uncover emerging trends. We observe a rise in vague research framings, raising concerns about the clarity of current framing practices in AFC research. We also find an increase in papers categorized under Scientific Curiosity, largely driven by benchmark studies examining LLM limitations. Lastly, we find a growing emphasis on generating fact-checking justifications to assist human fact-checkers, while fewer studies propose fully automating the process.
@@
Scientific inquiry Explores models or data without specific application goals. Scientific curiosity Scientific curiosity
@@
Table 1: Cross-domain research framings, with corresponding instantiations in AFC, HS. Clear framings (blue) link technical contributions to explicit purposes/users; vague framings (pink) lack specificity about use/users/goals.
@@ 2 Cross-domain research framing
To assess framing quality, we distinguish between clear and vague cases. Clear framings specify the artefact’s purpose, users, and goals; vague ones omit key components or introduce inconsistencies that obscure the system’s intended use. We note that theoretical research without an immediate ends or intended user is not the same as vagueness. If a paper clearly states that the goal is theoretical exploration or to answer some research question, that corresponds to Scientific inquiry, a clear framing. Vagueness arises when this intent is unstated, e.g., authors claiming that their theoretical findings aim to “fight hate” without discussing how.
@@
LLM safety represents a clear framing focused on preventing harmful LLM outputs through detoxification, filtering, or alignment. Vague data analysis refers to papers combining scientific inquiry with social aims (e.g.,“fighting hate”), without a clear link between analysis and real-world impact.
@@ 3.1 Step 1: Epistemic Element Extraction
Given a paragraph $p$ and an epistemic element $e\in E$, we define a set of candidate responses $R_{e}$ (including not specified). We frame the extraction as a QA task, using a language model $M$ to compute response likelihoods:
$s_{p,r_{e}}=\log P_{M}(r_{e}\mid p,q_{e})$ (1)
where $q_{e}$ is a predefined prompt targeting element $e$ (e.g., “Who are the individuals or entities that the authors identify as expected to act on the released artefacts?”) and $r_{e}\in R_{e}$ is a candidate label (e.g. journalists).
@@
We adapt Kuhn et al. (2023) to cluster sampled outputs $\{x_{1},\ldots,x_{k}\}$ into semantically equivalent classes $C=\{c_{1},\ldots,c_{m}\}$. Two outputs $x_{i},x_{j}$ are grouped together if they entail each other bidirectionally (i.e., NLI returns entail in both directions). Class likelihoods are computed via:
$s_{c_{i}}=\log\sum_{x\in c_{i}}\exp(s_{p,x})$ (2)
@@ 3.2 Step 2: Research Framing Ranking
In this step, we estimate the likelihood of each research framing $f_{j}$ by linking predicted epistemic elements through interpretable logical rules. Each research framing $f_{j}$ is defined by a boolean rule $R_{j}$ that specifies how combinations of epistemic elements must co-occur to support that framing. These rules are constructed from two sources: (1) a CART-based decision tree Breiman et al. (1984) trained on annotated data, and (2) manual refinement based on framing definitions (Appendix C.1). An example is shown in Figure 3 in the appendix.
To compute the likelihood of a research framing, we evaluate each logical rule $R_{j}$ over the predicted probabilities of its constituent epistemic elements. Conjunctions and disjunctions are computed as $\Pr(e_{1}\land e_{2})=\min\{\Pr(e_{1}),\Pr(e_{2})\}$ and
$\Pr(e_{1}\lor e_{2})=\max\{\Pr(e_{1}),\Pr(e_{2})\}$, respectively. This soft logic enables smooth aggregation of partial evidence across elements.
@@
Let a paper consist of paragraphs $\{p_{1},p_{2},\dots,p_{T}\}$, and let $\Pr(e_{k}\mid p_{t})$ denote the confidence score of epistemic element label $e_{k}$ in paragraph $p_{t}$. We compute the paper-level score for $e_{k}$ as $\Pr(e_{k})=\max_{t=1}^{T}\Pr(e_{k}\mid p_{t})$.
To model absence, we define the paper-level score for the not specified label as:
$\Pr(\textit{not specified}_{k})=\min_{t=1}^{T}\Pr(\textit{not specified}_{k}\mid p_{t})$
@@ 3.3 Research Framing Classification
To refine framing predictions, we use an LLM $M$ to incorporate document context and reason beyond deterministic rules. We first collect confidence scores from previous stages. Each framing $f_{j}$ is associated with a confidence score $\Pr(f_{j})$, and its supporting epistemic elements $E_{j}=\{e_{1},\dots,e_{n}\}$ have scores $\Pr(e_{i})$. We discretize each score into a confidence tier for each paper $j$:
$\text{Tier}(\Pr(z))$ =$=\begin{cases}\text{High}&\text{if }\Pr(z)\geq\tau_{h}\\ \text{Medium}&\text{if }\tau_{m}\leq\Pr(z)<\tau_{h}\\ \text{Low}&\text{if }\tau_{l}\leq\Pr(z)<\tau_{m}\end{cases}$
for all $\quad\text{for all }z\in\{f_{j}\}\cup E_{j}$
@@ 4 Research Framings in Hate Speech
The distribution of research framings in our annotated HS dataset reveals that Vague Identification is the most prevalent (44.74%). These papers present systems for detecting toxic content but fail to specify who will act on the outputs or how. Overall, vague research framings make up 64.47% of the dataset, underscoring a broader trend of ambiguity in the intended use of hate speech detection models (see Appendix D for distribution details). This is reinforced by the prevalence of Vague Data Analysis, where papers combine scientific inquiry with social aims (e.g., “fighting hate”) without articulating how the analysis supports real-world impact.
@@ 5.1 Implementation details
Absence Thresholding.
Papers often present multiple research framings across sections—e.g., one section may frame a tool as identifying misleading claims without specifying intended actors (Vague Identification), while another specifies its use by journalists. Strictly assigning $\infty$ to not specified labels when data actors appear in some paragraphs would incorrectly exclude Vague Identification, even if applicable elsewhere. To balance specificity with flexibility, we set a -3 threshold for not specified labels—low enough to avoid inflating irrelevant elements (as model-assigned likelihoods exceed -3) but high enough to retain research framings dependent on their absence.
@@
We define confidence tiers for both epistemic element and framing scores using the thresholds $\tau_{h}=-1$, $\tau_{m}=-2$, and $\tau_{l}=-3$, consistent with Gemini’s typical scoring range of approximately $[-3,0]$. These tiers are used to guide natural language justifications, which are passed to the LLM in conjunction with paper context, framing definitions, and in-context examples (see Figure 2). Using natural language labels for confidence levels improves the model’s reasoning by making the signal more interpretable.
@@ 5.2 Evaluation metrics
We evaluate performance for epistemic element extraction and research framing ranking using Filtered Mean Reciprocal Rank (fMRR; Bordes et al. (2013)), which averages the inverse rank of correct predictions while excluding higher-ranked known positives. That is, when multiple correct answers exist, we compute the rank of a target answer after removing other correct answers ranked above it. Also, some labels are unscored ($\infty$- ∞) if not predicted (e.g., professional journalists when only social media users are mentioned in the paper).
For scored labels, we compute standard fMRR. If a correct label is unscored, we estimate its rank as the number of scored labels plus half the unscored ones, assuming unscored items are randomly distributed. For example, with 2 scored and 10 unscored labels, the estimated rank is $72+\frac{10}{2}=$ 7, preventing artificially inflated results.
@@
$\text{rank}_{i}$ =$=\begin{cases}r_{i},&\text{if label }i\text{ is scored}\\ s+\frac{u}{2},&\text{if label }i\text{ is unscored}\end{cases}$ (4)
@@ 5.5 Research framing classification
Scientific inquiry Scientific curiosity 5 52.41 18.92 $\pm$ 4.26 20.24 $\pm$ 2.53 16 47.43 34.60 $\pm$ 3.62 48.84 $\pm$ 2.96
@@
To evaluate the impact of confidence levels in research framing ranking justifications on the LLM’s final classification, we analyze research framing retention rates. We find that 85.26% of research framings with high likelihoods (-1 to 0) are included in the final predictions, while 44.64% of medium-likelihood research framings are retained. In contrast, only 8.47% of the low-likelihood research framings are kept. This suggests that while the LLM considers confidence scores, it does not rely on them entirely and applies its own reasoning in the final classification. Finally, the system excels at detecting vague research framings, achieving F-scores of 58.91% (AFC) and 69.93% (HS) with precision rates of 76.16% and 86.24%.
@@ 6 Automated analysis of AFC literature
We also observe a rise in Vague Debunking framings and a decline in Vague Opposition. The former describes papers that propose automated debunking without clarifying whether models support or replace human fact-checkers; the latter presents a broad goal without specifying how model outputs are used to achieve it. While our system reliably detects vagueness, it sometimes defaults to Vague Debunking due to its broader scope—e.g., in Chamoun et al. (2023), it correctly identifies ambiguity in deployment but infers intended use based on model design, leading to misclassification.

### v2

@@ Abstract
Clarifying the research framing of NLP artefacts (e.g., models, datasets, etc.) is crucial to aligning research with practical applications when researchers claim that their findings have real-world impact. Recent studies manually analyzed NLP research across domains, showing that few papers explicitly identify key stakeholders, intended uses, or appropriate contexts. In this work, we propose to automate this analysis, developing a three-component system that infers research framings by first extracting key elements (means, ends, stakeholders), then linking them through interpretable rules and contextual reasoning. We evaluate our approach on two domains: automated fact-checking using an existing dataset, and hate speech detection for which we annotate a new dataset [^1] Code and annotations available at our GitHub repository. —achieving consistent improvements over strong LLM baselines. Finally, we apply our system to recent automated fact-checking papers and uncover three notable trends: a rise in underspecified research goals, increased emphasis on scientific exploration over application, and a shift toward supporting human fact-checkers rather than pursuing full automation.
@@ 1 Introduction
As NLP systems see wider real-world deployment, concerns over misuse and unintended consequences have intensified focus on responsible AI practices—highlighting the need for ethical integration, risk mitigation, and attention to societal impact (Noble, 2018; Bender et al., 2021; Birhane et al., 2022). A critical but often overlooked aspect of responsible NLP research is research framing—how a paper positions the purpose and value of its work by connecting what it does (means), why it does it (ends), and for whom it is relevant (stakeholders). In this context, framing highlights the methodological specification of an artefact’s role within the research ecosystem. This contrasts with political or media framing, which emphasizes persuasion in discourse, and semantic framing, which concerns meaning in linguistic structure Card et al. (2016).
When poorly articulated, research framings can obscure a system’s scope, hinder accountability, and increase the risk of misuse. Figure 1 illustrates this point by comparing two NLP papers: one clearly intends to support professional fact-checkers with specified means, ends, and users; the other, on hate speech detection, lacks any mention of who would use the model’s outputs. Such omissions risk misinterpretation—e.g., framing a label-only classification model as an automated decision tool overlooks the need for context and explanation, risking misuse in high-stakes settings.
@@
Figure 1: Side-by-side examples of research framings. The fact-checking paper (left, Konstantinovskiy et al. (2021)) presents a clear assisted external fact-checking framing with specified users, goals, and methods. The hate speech paper (right, Ousidhoum et al. (2019)) has an underspecified framing due to unspecified model users.
This concern has motivated several analyses of how NLP research is framed. Schlichtkrull et al. (2023) first formalized the notion of research framings (which they termed “epistemic narratives”) and introduced epistemic elements—textual signals that define a paper’s framing by specifying components such as means, ends and stakeholders. Analyzing 100 automated fact-checking papers, they found frequent misalignments between goals and methods, underspecified or missing stakeholder definitions, and weak links between these elements. Separately, Liu et al. (2023) reviewed 333 summarization papers and found that fewer than 15% meaningfully addressed responsible AI concerns.
@@
    Epistemic element extraction, i.e., identifying parts of the introductory text that specify the components of a paper’s research framing, such as its aims, means and stakeholders.
@@
We evaluate our approach (§5) in two domains: automated fact-checking (AFC) and hate speech detection (HS). For AFC, we use the dataset released by Schlichtkrull et al. (2023); for HS, we manually annotate 49 highly cited papers (§4) with epistemic elements, which we use to infer their research framings. Our approach outperforms a strong LLM baseline in both domains, offering more accurate research framing predictions. It also achieves high precision in identifying underspecified framings with misaligned or incomplete elements.
Finally, we apply our system (§6) to recent AFC papers to uncover emerging trends. We observe a rise in underspecified research framings, raising concerns about the clarity of current framing practices in AFC research. We also find an increase in papers categorized under Scientific Curiosity, largely driven by benchmark studies examining LLM limitations. Lastly, we find a growing emphasis on generating fact-checking justifications to assist human fact-checkers, while fewer studies propose fully automating the process.
@@
Knowledge exploration Explores models or data without specific application goals. Scientific curiosity Scientific curiosity
@@
Table 1: Cross-domain research framings, with corresponding instantiations in AFC, HS. Clear framings (blue) link technical contributions to explicit purposes/users; underspecified framings (pink) lack specificity about use/users/goals.
@@ 2 Cross-domain research framing
To assess framing quality, we distinguish between clear and underspecified cases. Clear framings specify the artefact’s purpose, users, and goals; underspecified ones omit key components or introduce inconsistencies that obscure the system’s intended use. We define underspecification relative to how clearly authors articulate and connect epistemic elements at the time of writing. In other words, it reflects presentation clarity in its contemporary context, not compliance with future or evolving community standards. This time-bounded definition makes the schema generalizable, since it provides a consistent basis for analyzing papers across different publication periods.
We also note that theoretical research without an immediate ends or intended user is not the same as underspecification. If a paper clearly states that the goal is theoretical exploration or to answer some research question, that corresponds to Knowledge exploration, a clear framing. Underspecification arises when this intent is unstated, e.g., authors claiming that their theoretical findings aim to “fight hate” without discussing how.
@@
LLM safety represents a clear framing focused on preventing harmful LLM outputs through detoxification, filtering, or alignment. Vague data analysis refers to papers combining knowledge exploration with social aims (e.g.,“fighting hate”), without a clear link between analysis and real-world impact.
@@ 3.1 Step 1: Epistemic Element Extraction
Given a paragraph $p_{i}$ from a selected paper and an epistemic element $E_{j}\in\mathcal{E}$ (e.g. data actors), we define the corresponding candidate label set $\mathcal{R}(E_{j})$ (including not specified). We frame the extraction as a QA task, using a language model $M$ to compute response likelihoods:
$s_{p_{i},r_{E_{j},k}}=\log P_{M}(r_{E_{j},k}\mid p_{i},q_{E_{j}})$ (1)
where $q_{E_{j}}$ is a predefined prompt targeting element $E_{j}$ (e.g., “Who are the individuals or entities that the authors identify as expected to act on the released artefacts?”) and $r_{E_{j},k}\in\mathcal{R}(E_{j})$ is a candidate label (e.g. journalists).
@@
We adapt Kuhn et al. (2023) to cluster sampled outputs $\{x_{1},\ldots,x_{n}\}$ into semantically equivalent classes $C=\{c_{1},\ldots,c_{m}\}$. Two outputs $x_{a},x_{b}$ are grouped together if they entail each other bidirectionally (i.e., NLI returns entail in both directions). Class likelihoods are computed via:
$s_{c_{j}}=\log\sum_{x\in c_{j}}\exp(s_{p_{i},x})$ (2)
@@ 3.2 Step 2: Research Framing Ranking
In this step, we estimate the likelihood of each research framing $f_{i}\in\mathcal{F}$ by linking predicted epistemic elements through interpretable logical rules. Each research framing $f_{i}$ is defined by a boolean rule $\mathcal{L}_{i}:\{r_{E_{1},k_{1}},r_{E_{2},k_{2}},\dots,r_{E_{m},k_{m}}\}\implies f_{i}$ that specifies how combinations of epistemic elements must co-occur to support that framing. These rules are constructed from two sources: (1) a CART-based decision tree Breiman et al. (1984) trained on annotated data, and (2) manual refinement based on framing definitions (Appendix C.1). An example is shown in Figure 3 in the appendix.
To compute the likelihood of a research framing, we evaluate each logical rule $\mathcal{L}_{i}$ over the predicted probabilities of its constituent epistemic elements. Conjunctions and disjunctions are computed as
$\Pr(r_{E_{a},k}\land r_{E_{b},l})=\min\{\Pr(r_{E_{a},k}),\Pr(r_{E_{b},l})\},$
$\Pr(r_{E_{a},k}\lor r_{E_{b},l})=\max\{\Pr(r_{E_{a},k}),\Pr(r_{E_{b},l})\}.$
. This soft logic enables smooth aggregation of partial evidence across elements.
@@
Let a paper consist of paragraphs $\{p_{1},p_{2},\dots,p_{T}\}$, and let $\Pr(r_{E_{j},k}\mid p_{t})$ denote the confidence score of label $r_{E_{j},k}$ for element $E_{j}$ in paragraph $p_{t}$. We compute the paper-level score as:
$\Pr(r_{E_{j},k})=\max_{t=1}^{T}\Pr(r_{E_{j},k}\mid p_{t}).$
To model absence, we define the paper-level score for the not specified label of element $E_{j}$ as:
$\Pr(r_{E_{j},\text{ns}})=\min_{t=1}^{T}\Pr(r_{E_{j},\text{ns}}\mid p_{t}).$
@@ 3.3 Research Framing Classification
To refine framing predictions, we use an LLM $M$ to incorporate document context and reason beyond deterministic rules. We first collect confidence scores from previous stages. Each framing $f_{i}$ is associated with a confidence score $\Pr(f_{i})$, and its supporting epistemic element labels $S_{i}$, $\quad\text{as specified by }\mathcal{L}_{i}$, have scores $\Pr(r_{E_{j},k})$. We discretize each score into a confidence tier for each paper:
$\text{Tier}(\Pr(z))$ $=\begin{cases}\text{High}&\text{if }\Pr(z)\geq\tau_{h}\\ \text{Medium}&\text{if }\tau_{m}\leq\Pr(z)<\tau_{h}\\ \text{Low}&\text{if }\tau_{l}\leq\Pr(z)<\tau_{m}\end{cases}$
$\quad\text{for all }z\in\{f_{i}\}\cup S_{i}$
@@ 4 Research Framings in Hate Speech
The distribution of research framings in our annotated HS dataset reveals that Vague Identification is the most prevalent (44.74%). These papers present systems for detecting toxic content but fail to specify who will act on the outputs or how. Overall, underspecified research framings make up 64.47% of the dataset, underscoring a broader trend of ambiguity in the intended use of hate speech detection models (see Appendix D for distribution details). This is reinforced by the prevalence of Vague Data Analysis, where papers combine knowledge exploration with social aims (e.g., “fighting hate”) without articulating how the analysis supports real-world impact.
@@ 5.1 Implementation details
Absence Thresholding
Papers often present multiple research framings across sections—e.g., one section may frame a tool as identifying misleading claims without specifying intended actors (Vague Identification), while another specifies its use by journalists. Strictly assigning -$\infty$ to not specified labels when data actors appear in some paragraphs would incorrectly exclude Vague Identification, even if applicable elsewhere. To balance specificity with flexibility, we set a -3 threshold for not specified labels—low enough to avoid inflating irrelevant elements (as model-assigned likelihoods exceed -3) but high enough to retain research framings dependent on their absence.
@@
We define confidence tiers for both epistemic element and framing scores using the thresholds $\tau_{h}=-1$, $\tau_{m}=-2$, and $\tau_{l}=-3$, consistent with Gemini’s typical scoring range of approximately [3,0][-3,0]. These tiers are used to guide natural language justifications, which are passed to the LLM in conjunction with paper context, framing definitions, and in-context examples (see Figure 2). Using natural language labels for confidence levels improves the model’s reasoning by making the signal more interpretable.
@@ 5.2 Evaluation metrics
We evaluate performance for epistemic element extraction and research framing ranking using Filtered Mean Reciprocal Rank (fMRR; Bordes et al. (2013)), which averages the inverse rank of correct predictions while excluding higher-ranked known positives. That is, when multiple correct answers exist, we compute the rank of a target answer after removing other correct answers ranked above it. Also, some labels are unscored (-$\infty$) if not predicted (e.g., professional journalists when only social media users are mentioned in the paper).
For scored labels, we compute standard fMRR. If a correct label is unscored, we estimate its rank as the number of scored labels plus half the unscored ones, assuming unscored items are randomly distributed. For example, with 2 scored and 10 unscored labels, the estimated rank is $2+\frac{10}{2}=7$, preventing artificially inflated results.
@@
[…]
@@ 5.5 Research framing classification
Knowledge exploration Scientific curiosity 5 52.41 18.92 $\pm$ 4.26 20.24 $\pm$ 2.53 16 47.43 34.60 $\pm$ 3.62 48.84 $\pm$ 2.96
@@
To evaluate the impact of confidence levels in research framing ranking justifications on the LLM’s final classification, we analyze research framing retention rates. We find that 85.26% of research framings with high likelihoods (-1 to 0) are included in the final predictions, while 44.64% of medium-likelihood research framings are retained. In contrast, only 8.47% of the low-likelihood research framings are kept. This suggests that while the LLM considers confidence scores, it does not rely on them entirely and applies its own reasoning in the final classification. Finally, the system excels at detecting underspecified research framings, achieving F-scores of 58.91% (AFC) and 69.93% (HS) with precision rates of 76.16% and 86.24%.
@@ 6 Automated analysis of AFC literature
We also observe a rise in Vague Debunking framings and a decline in Vague Opposition. The former describes papers that propose automated debunking without clarifying whether models support or replace human fact-checkers; the latter presents a broad goal without specifying how model outputs are used to achieve it. While our system reliably detects underspecification, it sometimes defaults to Vague Debunking due to its broader scope—e.g., in Chamoun et al. (2023), it correctly identifies ambiguity in deployment but infers intended use based on model design, leading to misclassification.